OpenStudy (anonymous):

The area of a square is 625 cm^2. What is the measure of the diagonal to the nearest tenth?

OpenStudy (anonymous):

if the area of a square is 625 cm2, what is its side?

OpenStudy (anonymous):

It doesn't say

OpenStudy (anonymous):

The Sides, since its a square, are all the same. Thus The length, a, and the width,b, are equal to each other a=b. Then you utilize the Pythagorean Theorem, \[a^{2}+b ^{2}=c ^{2}\] Where a and b are the sides and c is the hypotenuse. Sides are, \[625cm^2 = a^2, a=\sqrt{625}cm\] \[\sqrt{625}^2 + \sqrt{625}^2 = c^2 \] C = 35.36